# Structure Formation with Dark Magnetohydrodynamics

### v1

@@ Abstract
Long-range interactions in the dark sector can give rise to collective plasma phenomena that are capable of modifying the evolution of dark matter halos. We present the first study of gravitational collapse in a secluded dark $U(1)_{D}$ model using a magnetohydrodynamic description of the dark matter. We show that dark magnetic fields generate an anisotropic pressure that alters the Jeans scale and suppresses small-scale power in a direction-dependent manner. For a range of primordial magnetic spectral indices, this effect produces distinctive modifications to the linear matter power spectrum. We find that current observations cannot yet constrain viable dark magnetic fields, as CMB tensor modes mostly provide more stringent constraints. Nevertheless, forthcoming high-resolution probes of the matter power spectrum (CMB-HD lensing, HERA, and EDGES) will be able to test these predictions and are sensitive to dark charge-to-mass ratios in the range $10^{-20}\,\text{GeV}^{-1}\lesssim q_{\chi}/m_{\chi}\lesssim 10^{-14}\,\text{GeV}^{-1}$.
@@ I Introduction
A majority of work in SIDM has focused on the scenario in which DM particles interact through a short-range force mediated by a boson with mass much greater than the typical momentum transfer, $q\sim v_{\rm DM}m_{\rm DM}$, where $v_{\rm DM}$ and $m_{\rm DM}$ are the typical velocity and mass of DM particles. In this limit, particles only interact via individual hard scattering events. Recently, there has been growing interest in the case in which DM possesses coherent, long-range self-interactions Bogorad:2023wzn. A subset of these models consider the scenario in which the dark force is mediated by an ultralight vector Ackerman_2009; Lasenby:2020; Cruz_2023; Giffin; Medvedev:2024kjh; Khlopov_1; Khlopov_2; Khlopov_3; Ackerman:2008kmp; Agrawal:2016quu; Cyr-Racine:2012tfp. If the particle collision rate is sufficiently small, these long-range interactions are dominated by collective effects. For instance, DM may be susceptible to velocity-space instabilities that lead to resonant exchange of energy between particles and (mediator) waves and can dramatically alter DM kinematics (see, e.g., Bhattacharjee2017 for a review of such processes in Standard Model plasmas).
In such scenarios, formation of plasma instabilities may occur in the dark sector, dramatically altering the system’s dynamics. The possibility of dark-sector plasma instabilities has been explored in several contexts Ackerman_2009; Lasenby:2020; Cruz_2023. In the visible sector, analyses of linear instabilities in environments such as supernova remnants, the CMB, and merging galaxy clusters Li_2020; Cruz_2023 have yielded stringent constraints on models where dark-sector particles interact with Standard Model (SM) electrons, since electron–electron collisions efficiently quench instability growth. In contrast, constraints on secluded dark sectors, where such collisional damping is absent, are considerably weaker. In these models, both electromagnetic and electrostatic modes in the dark plasma can grow largely unimpeded, allowing instabilities to persist over cosmological timescales. This broader instability landscape motivates a detailed study of their impact on structure formation and on astrophysical observables.
Beyond linear theory, numerical simulations have begun to explore the nonlinear dynamics of dark-sector plasmas. Hydrodynamic simulations of cluster collisions under a two-component DM model, treating part of the DM as a fluid-like plasma, have been able to reproduce density maps that closely resemble Bullet Cluster and Abell 520 in the case where a significant fraction of the DM behaves as a plasma Heikinheimo_2015; Spethmann_2017; Heikinheimo_2018. More recently, fully kinetic simulations of counter-streaming dark-matter flows, combined with observations of the Bullet Cluster, have further constrained the parameter space of secluded dark $U(1)_{D}$ Giffin. These studies highlight the potentially significant astrophysical consequences of plasma-like interactions in the dark sector.
@@
    We begin in section II with a brief review of the Classical Gravitational Instability without any alterations from dark sector plasma effects. This is a primer for the work done in section III.
@@
    In section III we introduce the dark $U(1)_{D}$ model that we consider throughout this work, and couple dark sector plasma effects to the Classical Jeans Instability, demonstrating a directionally dependent alteration to both the Jeans Length and group velocity in the presence of an external dark magnetic field.
@@
    We then present the alterations to the linear matter power spectrum under this model in section V. Scanning over parameter space and performing a $\chi^{2}$ analysis against current and future measurements of the primordial matter power spectrum, we also show our constraints on the background dark magnetic field, as well as the charge-to-mass ratio for this model for various different spectral indices for the primordial magnetic field spectrum.
@@
    Finally, with section VI we conclude the paper with a discussion of not only the constraints we derived, but also possible future extensions of our work, such as the consideration of kinetic mixing with the Standard Model photon and the possibility of halo triaxiality as a result of anisotropic collapse due to the directionally-dependent alteration to the Jeans Length.
@@ II Gravitational Collapse and Classical Jeans Length
Here, $c_{s}$ denotes the speed of sound in the medium, and $\rho_{0}$ is the unperturbed mass density of the ensemble. This defines the critical length at which the gas becomes stable against gravitational collapse. Specifically, only perturbations with wavelengths that exceed the Jeans length $\lambda>\lambda_{J}$ will trigger instabilities binney_galactic_2008.
@@ III Evolution of Dark Magnetized Plasma
To allow for collective plasma effects, we consider a dark sector consisting of dark positrons and electrons of equal abundance that interact through a massless $U(1)_{D}$ gauge boson, which we will refer to as a “dark photon”. The dark sector Lagrangian for this model takes the form
@@
where $\chi$, $\bar{\chi}$ are dark electrons and positrons, respectively, with charge $q_{\chi}$ and mass $m_{\chi}$, and $A^{\prime}$ is the dark photon. An additional term of the form $\frac{\epsilon}{2}F^{\prime\mu\nu}F_{\mu\nu}$ would allow for the possibility of kinetic mixing between the dark photon and the Standard Model photon. In this work, we will only consider the case where kinetic mixing does not occur, giving rise to a secluded dark sector. Note that if $\epsilon$ is set to zero at some high energy scale, it will remain zero under renormalization group evolution unless there are heavy states charged under both $U(1)_{D}$ and $U(1)_{\rm EM}$ Ackerman:2008kmp. We leave the study of $\epsilon\neq 0$ for future work and restrict the model to the scenario of a massless dark photon.
Previous works have utilized the evolution of beam instabilities in plasmas to derive constraints on such a model Cruz_2023; Li_2020; Giffin. For the purpose of characterizing the effects of plasma instabilities and how they alter structure formation in the universe, beam-type instabilities do not provide the correct picture for this scenario. Instead, we consider a homogeneous and isotropic, net-neutral plasma of particles to characterize the criteria for gravitational collapse in the presence of collective plasma effects.
The evolution of a plasma can be well-approximated by a fluid model. Though this assumption is typically made in highly collisional systems, it can also provide a framework to study the evolution of collisionless systems whose velocity distributions remain approximately Maxwellian. In a collisional plasma, particle collisions drive the distribution function towards a Maxwellian, as explained by Boltzmann’s H-theorem. Even in collisionless plasmas, collective kinetic processes such as phase mixing, Landau damping, wave-particle scattering, and microinstabilities can produce distributions that are close to Maxwellian over macroscopic scales, despite the absence of true collisions. These systems may, however, evolve away from Maxwellians (e.g., toward power-laws) on longer timescales Ewart:2024bnh. We leave a more detailed treatment of these effects to future work. Treating the plasma as a fluid yields the system of partial differential equations often referred to as the magnetohydrodynamic (MHD) equations. Adopting a system of units in which $\hbar=c=\epsilon_{0}=\mu_{0}=1$, the system of equations is
@@
Here $\vec{U}$ is the bulk velocity of the DM fluid, $\rho_{c}$ and $\rho_{m}$ are the charge and mass density of the fluid, respectively. $\stackrel{{\scriptstyle\leftrightarrow}}{{P}}$ is the pressure tensor containing only diagonal elements of $p_{\perp}$ in the directions perpendicular to the magnetic field and $p_{\parallel}$ in the direction parallel to the magnetic field. Additionally, we have assumed the Chew-Goldberger-Low (CGL) equation of state for cylindrically symmetric magnetized plasmas CGL:1956. The last equation is a specific scenario of the generalized Ohm’s law discussed in Appendix A.
@@ III.2 Perpendicular Propagation
These modes only contain a single solution for $\omega$:
@@ IV.1 Dark Magnetic Field Evolution
where $\vec{B}_{\rm{com}}=a^{2}\vec{B}_{\rm phys}$ is the comoving magnetic field, $\vec{B}_{\rm phys}$ is the physical field, and $\vec{v_{b}}$ is the bulk velocity of the plasma. On large comoving scales, $\ell\gg v_{b}/aH$, the comoving magnetic field remains constant, giving rise to the standard evolution of the physical field, $\vec{B}_{\rm phys}\propto a^{-2}$. Below this scale, MHD turbulence develops, suppressing both density perturbations and magnetic fields Banerjee:2004df; Trivedi:2018ejz; Jedamzik:2018itu. The transition scale is
@@
where $B_{i}(\vec{k})\equiv\int d^{3}xB(\vec{x})\exp(i\vec{k}\cdot\vec{x})$ is the Fourier transform of the magnetic field. The power spectrum is typically parameterized as follows:
@@
where $n$ is the model-dependent spectral tilt, and $\ell$ is the transition scale in (27). It is assumed that MHD dissiption is the only cutoff in the PMF power spectrum. PMFs generated during inflation result in nearly scale-invariant spectra with $n\approx-3$. PMFs generated after inflation (e.g., in a phase transition) typically have blue spectra with n>3n>-3. Constraints on the magnetic field are typically quoted in terms of the comoving field strength, $B_{\rm com}$, at some reference scale $\lambda_{B}$, typically 1 Mpc:
$B_{\rm com}(\lambda)\equiv\left\langle B^{2}(\lambda)\right\rangle^{1/2}\propto\left[\displaystyle\int_{0}^{\infty}k^{2}P_{B}(k)dk\right]^{1/2}=B(\lambda_{B})\left(\frac{\lambda}{\lambda_{B}}\right)^{-\frac{n+3}{2}}\Gamma^{1/2}\left[(n+3)/2\right],$ (30)
@@
In the following subsection, we summarize current constraints on SM PMFs and show how these bounds can be mapped onto constraints on dark PMFs.
@@ IV.2 Constraints on Standard Model Primordial Magnetic Fields
A common mechanism for primordial seed magnetic field generation is the so-called Biermann battery, in which the plasma pressure gradient and temperature gradient are not aligned 1962ApJ...136..615M; 1950ZNatA...5...65B. In SM plasmas, the misalignment arises from electron-ion drift effects originating from the mass hierarchy between the two charged species. The Biermann battery is inoperative for symmetric charge species, such as in the dark $U(1)_{D}$ model under consideration. Kinetic plasma instabilities, such as the Weibel instability PhysRevLett.2.83; 10.1063/1.1705933 can also generate and amplify magnetic fields in plasmas with anisotropic velocity distributions Medvedev_2006; 2009ApJ...693.1133L. In the SM sector, this is typically studied in the context of merging and collapsing galaxies. These mechanisms could be extended to the dark sector as well if significant velocity anisotropies can be created.
@@
Alternatively, more exotic mechanisms have been considered to generate seed primordial magnetic fields. During the inflation, primordial magnetic fields can be generated by explicitly breaking the conformal invariance of the electromagnetic action. Refs. PhysRevD.37.2743; 1992ApJ...391L...1R; PhysRevD.48.2499; PhysRevLett.75.3796; PhysRevD.62.067301; PhysRevD.89.063002; Fujita_2015; KANDUS20111; Subramanian_2010; Subramanian_2016; Martin_2008 consider direct couplings of the electromagnetic fields to other fields that explicitly break the conformal symmetry during inflation. Additionally, models have been suggested that add an $RA_{\mu}A^{\mu}$ term to the action but are strongly disfavored since they contain ghosts Himmetoglu_2009.
@@
Lastly, phase transitions in the early universe could lead to the generation of magnetic fields. In the SM sector, electroweak phase transition and the QCD transition are often considered Brandenburg_1996; Banerjee_2004. The dark $U(1)_{D}$ could be the result of some other phase transition in the dark sector or from couplings to SM. First-order phase transitions provide ideal generation mechanisms as colliding bubble walls can seed turbulent fields. Ref. VACHASPATI1991258 also claims that the gradient of the Higgs vacuum expectation value can induce electromagnetic fields at an efficient rate during electroweak phase transition.
@@
After seed field generation, the primordial magnetic fields may be enhanced exponentially by dynamo processes. For turbulent fields, the fluctuation dynamo 1968JETP...26.1031K provides the most promising growth mechanism. Given we are in a regime of very high magnetic Reynolds number, the seed field can grow rapidly 1986ESASP.251..557K but may not grow optimally due to the DM’s compressible nature 1985ZhETF..88..487K.
@@
Any magnetic field (SM or dark) will contribute to the Hubble parameter governing the expansion rate of the universe. If magnetic fields are too large, then the correct abundance of 4He will not be produced. Ref. Grasso_1996 finds that an upper limit of $B_{\rm{phys}}<2\times 10^{11}$ Gauss can be placed at $T=10^{9}$ K for magnetic fields that are uniform over scales larger than the Hubble radius. This corresponds to $B_{\text{phys}}<0.7\,\mu\text{G}$ today.
Random magnetic fields may also generate tensor perturbations leading to anisotropies in the CMB through their coupling to gravitational waves. Ref. PhysRevD.61.043001 find that for a scale invariant spectrum on galactic scales, $B_{\text{phys}}<1$ nG. For a spectral index n<3/2n<-3/2, $B_{\lambda}<7.9e^{3n}\,\mu\text{G}$ and for a spectral index of n>3/2n>-3/2, $B_{\lambda}<95e^{0.37n}\,\text{nG}$ where $B_{\lambda}$ is the amplitude at $\lambda=0.1\,h^{-1}\,\rm{Mpc}$. These constraints, translated to constraints at $z=30$ using Eq. (31), are shown in Fig. 1.
@@
Figure 1: Constraints on the magnetic field strength at different comoving scales: $\lambda=0.1$ Mpc (dark blue), 1 Mpc (medium blue), and 10 Mpc (light blue) at redshift $z=30$ as a function of spectral tilt of the magnetic power spectrum, $n$, reproduced from Ref. PhysRevD.61.043001.

### v2

@@ Abstract
Long-range interactions in the dark sector can give rise to collective plasma phenomena that are capable of modifying the evolution of dark matter halos. We present the first study of gravitational collapse in a secluded dark $U(1)_{D}$ model using a magnetohydrodynamic description of the dark matter. We show that dark magnetic fields generate an anisotropic pressure that alters the Jeans scale and suppresses small-scale power in a direction-dependent manner. For a range of primordial magnetic spectral indices, this effect produces distinctive modifications to the linear matter power spectrum. We find that current observations cannot yet constrain viable dark magnetic fields, as cosmic microwave background (CMB) tensor modes mostly provide more stringent constraints. Nevertheless, forthcoming high-resolution probes of the matter power spectrum (CMB-HD lensing, HERA, and EDGES) will be able to test these predictions and are sensitive to dark charge-to-mass ratios in the range $10^{-20}\,\text{GeV}^{-1}\lesssim q_{\chi}/m_{\chi}\lesssim 10^{-14}\,\text{GeV}^{-1}$.
@@ I Introduction
A majority of work in SIDM has focused on the scenario in which DM particles interact through a short-range force mediated by a boson with mass much greater than the typical momentum transfer, $q\sim v_{\rm DM}m_{\rm DM}$, where $v_{\rm DM}$ and $m_{\rm DM}$ are the typical velocity and mass of DM particles. In this limit, particles only interact via individual hard scattering events. Recently, there has been growing interest in the case in which DM possesses coherent, long-range self-interactions [17]. A subset of these models consider the scenario in which the dark force is mediated by an ultralight vector [2, 60, 24, 31, 68, 15, 16, 56, 3, 7, 25]. If the particle collision rate is sufficiently small, these long-range interactions are dominated by collective effects. For instance, DM may be susceptible to velocity-space instabilities that lead to resonant exchange of energy between particles and (mediator) waves and can dramatically alter DM kinematics (see, e.g., [46] for a review of such processes in Standard Model plasmas).
In such scenarios, formation of plasma instabilities may occur in the dark sector, dramatically altering the system’s dynamics. The possibility of dark-sector plasma instabilities has been explored in several contexts [2, 60, 24]. In the visible sector, analyses of linear instabilities in environments such as supernova remnants, the CMB, and merging galaxy clusters [63, 24] have yielded stringent constraints on models where dark-sector particles interact with Standard Model (SM) electrons, since electron–electron collisions efficiently quench instability growth. In contrast, constraints on secluded dark sectors, where such collisional damping is absent, are considerably weaker. In these models, both electromagnetic and electrostatic modes in the dark plasma can grow largely unimpeded, allowing instabilities to persist over cosmological timescales. This broader instability landscape motivates a detailed study of their impact on structure formation and on astrophysical observables.
Beyond linear theory, numerical simulations have begun to explore the nonlinear dynamics of dark-sector plasmas. Hydrodynamic simulations of cluster collisions under a two-component DM model, treating part of the DM as a fluidlike plasma, have been able to reproduce density maps that closely resemble Bullet Cluster and Abell 520 in the case where a significant fraction of the DM behaves as a plasma [47, 85, 48]. More recently, fully kinetic simulations of counter-streaming dark-matter flows, combined with observations of the Bullet Cluster, have further constrained the parameter space of secluded dark $U(1)_{D}$ [31]. These studies highlight the potentially significant astrophysical consequences of plasmalike interactions in the dark sector.
@@
    We begin in Sec. II with a brief review of the classical gravitational instability without any alterations from dark sector plasma effects. This is a primer for the work done in Sec. III.
@@
    In Sec. III we introduce the dark $U(1)_{D}$ model that we consider throughout this work, and couple dark sector plasma effects to the classical Jeans instability, demonstrating a directionally dependent alteration to both the Jeans length and group velocity in the presence of an external dark magnetic field.
@@
    We then present the alterations to the linear matter power spectrum under this model in Sec. V. Scanning over parameter space and performing a $\chi^{2}$ analysis against current and future measurements of the primordial matter power spectrum, we also show our constraints on the background dark magnetic field, as well as the charge-to-mass ratio for this model for various different spectral indices for the primordial magnetic field spectrum.
@@
    Finally, with Sec. VI we conclude the paper with a discussion of not only the constraints we derived, but also possible future extensions of our work, such as the consideration of kinetic mixing with the Standard Model photon and the possibility of halo triaxiality as a result of anisotropic collapse due to the directionally dependent alteration to the Jeans length.
@@ II Gravitational Collapse and classical Jeans length
Here, $c_{s}$ denotes the speed of sound in the medium, and $\rho_{0}$ is the unperturbed mass density of the ensemble. This defines the critical length at which the gas becomes stable against gravitational collapse. Specifically, only perturbations with wavelengths that exceed the Jeans length $\lambda>\lambda_{J}$ will trigger instabilities [14].
@@ III Evolution of Dark Magnetized Plasma
To allow for collective plasma effects, we consider a dark sector consisting of dark positrons and electrons of equal abundance that interact through a massless $U(1)_{D}$ gauge boson, which we will refer to as a “dark photon.” The dark sector Lagrangian for this model takes the form
@@
where $\chi$, $\bar{\chi}$ are dark electrons and positrons, respectively, with charge $q_{\chi}$ and mass $m_{\chi}$, and $A^{\prime}$ is the dark photon. An additional term of the form $\frac{\epsilon}{2}F^{\prime\mu\nu}F_{\mu\nu}$ would allow for the possibility of kinetic mixing between the dark photon and the Standard Model photon. In this work, we will only consider the case where kinetic mixing does not occur, giving rise to a secluded dark sector. Note that if $\epsilon$ is set to zero at some high energy scale, it will remain zero under renormalization group evolution unless there are heavy states charged under both $U(1)_{D}$ and $U(1)_{\rm EM}$ [3]. We leave the study of $\epsilon\neq 0$ for future work and restrict the model to the scenario of a massless dark photon.
Previous works have utilized the evolution of beam instabilities in plasmas to derive constraints on such a model [24, 63, 31]. For the purpose of characterizing the effects of plasma instabilities and how they alter structure formation in the Universe, beam-type instabilities do not provide the correct picture for this scenario. Instead, we consider a homogeneous and isotropic, net-neutral plasma of particles to characterize the criteria for gravitational collapse in the presence of collective plasma effects.
The evolution of a plasma can be well-approximated by a fluid model. Though this assumption is typically made in highly collisional systems, it can also provide a framework to study the evolution of collisionless systems whose velocity distributions remain approximately Maxwellian. In a collisional plasma, particle collisions drive the distribution function toward a Maxwellian, as explained by Boltzmann’s H-theorem. Even in collisionless plasmas, collective kinetic processes such as phase mixing, Landau damping, wave-particle scattering, and microinstabilities can produce distributions that are close to Maxwellian over macroscopic scales, despite the absence of true collisions. These systems may, however, evolve away from Maxwellians (e.g., toward power-laws) on longer timescales [38]. We leave a more detailed treatment of these effects to future work. Treating the plasma as a fluid yields the system of partial differential equations often referred to as the magnetohydrodynamic (MHD) equations. Adopting a system of units in which $\hbar=c=\epsilon_{0}=\mu_{0}=1$, the system of equations is
@@
Here $\vec{U}$ is the bulk velocity of the DM fluid, $\rho_{c}$ and $\rho_{m}$ are the charge and mass density of the fluid, respectively. $\stackrel{{\scriptstyle\leftrightarrow}}{{P}}$ is the pressure tensor containing only diagonal elements of $p_{\perp}$ in the directions perpendicular to the magnetic field and $p_{\parallel}$ in the direction parallel to the magnetic field. Additionally, we have assumed the Chew-Goldberger-Low (CGL) equation of state for cylindrically symmetric magnetized plasmas [22]. The last equation is a specific scenario of the generalized Ohm’s law discussed in Appendix A.
@@ III.2 Perpendicular propagation
These modes only contain a single solution for $\omega$
@@ IV.1 Dark magnetic field evolution
where $\vec{B}_{\rm{com}}=a^{2}\vec{B}_{\rm phys}$ is the comoving magnetic field, $\vec{B}_{\rm phys}$ is the physical field, and $\vec{v_{b}}$ is the bulk velocity of the plasma. On large comoving scales, $\ell\gg v_{b}/aH$, the comoving magnetic field remains constant, giving rise to the standard evolution of the physical field, $\vec{B}_{\rm phys}\propto a^{-2}$. Below this scale, MHD turbulence develops, suppressing both density perturbations and magnetic fields [11, 90, 50, 57, 88]. The transition scale is
@@
where $B_{i}(\vec{k})\equiv\int d^{3}xB(\vec{x})\exp(i\vec{k}\cdot\vec{x})$ is the Fourier transform of the magnetic field. The power spectrum is typically parametrized as follows:
@@
where $n$ is the model-dependent spectral tilt, and $\ell$ is the transition scale in (27). It is assumed that MHD dissiption is the only cutoff in the PMF power spectrum. PMFs generated during inflation result in nearly scale-invariant spectra with $n\approx-3$. PMFs generated after inflation must be causally connected which constraints $n=2$ [35]. Constraints on the magnetic field are typically quoted in terms of the comoving field strength, $B_{\rm com}$, at some reference scale $\lambda_{B}$, typically 1 Mpc
$B_{\rm com}(\lambda)\equiv\left\langle B^{2}(\lambda)\right\rangle^{1/2}\propto\left[\displaystyle\int_{0}^{\infty}k^{2}P_{B}(k)dk\right]^{1/2}$
$=B(\lambda_{B})\left(\frac{\lambda}{\lambda_{B}}\right)^{-\frac{n+3}{2}}\Gamma^{1/2}\left[(n+3)/2\right],$ (30)
@@
Throughout the paper, we do not assume any particular mechanism to generate PMFs. However, in order to avoid the development of turbulence on scales of interest, we assume that the energy injection scale is much smaller than the transition scale in Eq. (27). In the following subsection, we summarize current constraints on SM PMFs and show how these bounds can be mapped onto constraints on dark PMFs.
@@ IV.2 Constraints on standard model primordial Magnetic Fields
A common mechanism for primordial seed magnetic field generation is the so-called Biermann battery, in which the plasma pressure gradient and temperature gradient are not aligned [70, 13]. In SM plasmas, the misalignment arises from electron-ion drift effects originating from the mass hierarchy between the two charged species. The Biermann battery is inoperative for charged species with equal masses, such as in the dark $U(1)_{D}$ model under consideration. In this instance, the contributions to the pressure gradient from the positive and negative species cancel. For more details see Appendix A. Kinetic plasma instabilities, such as the Weibel instability [94, 39] can also generate and amplify magnetic fields in plasmas with anisotropic velocity distributions [69, 61]. In the SM sector, this is typically studied in the context of merging and collapsing galaxies. These mechanisms could be extended to the dark sector as well if significant velocity anisotropies can be created.
@@
Alternatively, more exotic mechanisms have been considered to generate seed primordial magnetic fields. During the inflation, primordial magnetic fields can be generated by explicitly breaking the conformal invariance of the electromagnetic action. References [92, 78, 33, 41, 42, 9, 40, 51, 87, 89, 66] consider direct couplings of the electromagnetic fields to other fields that explicitly break the conformal symmetry during inflation. Additionally, models have been suggested that add an $RA_{\mu}A^{\mu}$ term to the action but are strongly disfavored since they contain ghosts [49].
@@
Lastly, phase transitions in the early Universe could lead to the generation of magnetic fields. In the SM sector, electroweak phase transition and the QCD transition are often considered [19, 10]. The dark $U(1)_{D}$ could be the result of some other phase transition in the dark sector or from couplings to SM. First-order phase transitions provide ideal generation mechanisms as colliding bubble walls can seed turbulent fields. Reference [93] also claims that the gradient of the Higgs vacuum expectation value can induce electromagnetic fields at an efficient rate during electroweak phase transition.
@@
After seed field generation, the primordial magnetic fields may be enhanced exponentially by dynamo processes. For turbulent fields, the fluctuation dynamo [54] provides the most promising growth mechanism. Given we are in a regime of very high magnetic Reynolds number, the seed field can grow rapidly [59] but may not grow optimally due to the DM’s compressible nature [53].
@@
Any magnetic field (SM or dark) will contribute to the Hubble parameter governing the expansion rate of the Universe. If magnetic fields are too large, then the correct abundance of 4He will not be produced. Ref. [44] finds that an upper limit of $B_{\rm{phys}}<2\times 10^{11}$ Gauss can be placed at $T=10^{9}$ K for magnetic fields that are uniform over scales larger than the Hubble radius. This corresponds to $B_{\text{phys}}<0.7\,\mu\text{G}$ today.
@@
Figure 1: Constraints on the magnetic field strength at different comoving scales: $\lambda=0.1$ Mpc (dark blue), 1 Mpc (medium blue), and 10 Mpc (light blue) at redshift $z=30$ as a function of spectral tilt of the magnetic power spectrum, $n$, reproduced from Ref. [34].
Random magnetic fields may also generate tensor perturbations leading to anisotropies in the CMB through their coupling to gravitational waves. Ref. [34] find that for a scale invariant spectrum on galactic scales, $B_{\text{phys}}<1$ nG. For a spectral index n<3/2n<-3/2, $B_{\lambda}<7.9e^{3n}\,\mu\text{G}$ and for a spectral index of n>3/2n>-3/2, $B_{\lambda}<95e^{0.37n}\,\text{nG}$ where $B_{\lambda}$ is the amplitude at $\lambda=0.1\,h^{-1}\,\rm{Mpc}$. These constraints, translated to constraints at $z=30$ using Eq. (31), are shown in Fig. 1.